OpenStudy (anonymous):

Evaluate the given expressions. All numbers are approximate. 15.66^2-(-4.017)^4/1.044(-3.68)

OpenStudy (anonymous):

you can use calculator for this problem, i think so.

OpenStudy (jack1):

what do u mean by evaluate...? solve?

OpenStudy (anonymous):

It just says to evaluate the expression and then solve them but keep in mind they are approximate

OpenStudy (anonymous):

For some reason I'm not getting the right answer

OpenStudy (jack1):

what answer are you getting?

OpenStudy (anonymous):

I got188.59

OpenStudy (jack1):

ok, I'm getting a different one, so lets take it line by line then

OpenStudy (anonymous):

There is a certain way this has to be done but I don't know

OpenStudy (jack1):

top line : 15.66^2-(-4.017)^4 so start with the powers 15.66^2 = 245.2356 (-4.017)^4 = 260.3798 so 15.66^2-(-4.017)^4 = 245.2356 - 260.3798 = -15.14 so numerator reduces to -15.14 all good so far?

OpenStudy (anonymous):

Oh okay I see

OpenStudy (jack1):

cools, now bottom line (denominator) 1.044(-3.68) = 1.044 times (-3.68) = -3.842

OpenStudy (jack1):

so your final equation (which you can pop straight into a calculator) is -15.14 / -3.842 = 15.14 / 3.842 =... ? @Ash90 ?

OpenStudy (anonymous):

Yes I have the bottom correct -3.84192

OpenStudy (jack1):

damn, no i had it wrong

OpenStudy (anonymous):

So my final answer should be 3.941

OpenStudy (jack1):

bottom = -3.84192 is correct

OpenStudy (jack1):

but the equation i think should be more like:

OpenStudy (jack1):

\[15.66^2-\frac{(-4.017)^4}{1.044 \times (-3.68)}\]

OpenStudy (jack1):

as there are no brackets around the 15 and the -4

OpenStudy (jack1):

sorry about that so start with the fraction (-4.017)^4 = 260.3798 over 1.044 times (-3.68) = -3.842 so 260.3798/ -3.842 = -67.77 therefore: (remembering 15.66^2 = 245.2356) 15.66^2 - (-67.77) = 245.2356 - (-67.77) = 245.2356 + 67.77 = ...?

OpenStudy (jack1):

= the final answer ;D @Ash90

OpenStudy (anonymous):

Okay thank you